# Tensor-based Graph Learning with Consistency and Specificity for Multi-view Clustering

### v2

@@ III-C Optimization
Let $e^{(v)}_{ij}=d_{v}-\mathrm{tr}(\mathbf{x}^{(v)}_{i}(\mathbf{x}^{(v)}_{j})^{T}$, we can further reformulate the problem in (LABEL:eq017) separately for each $i$
@@
Given that the $l_{2,1}$-norm of a tensor is defined as the sum of the $l_{2}$-norms for each fiber along the third mode, thus we have $\lVert\mathbf{W}_{(3)}\rVert_{2,1}=\lVert\mathcal{W}\rVert_{2,1}$. Similarly, according the definition of the $F$-norm of a tensor, it is straightforward to see $\lVert\mathbf{W}_{(3)}-(\mathbf{E}_{(3)}+\frac{\mathbf{Q}_{3_{(3)}}}{\rho})\rVert^{2}_{F}=\lVert\mathcal{W}-(\mathcal{E}+\frac{\mathcal{Q}_{3}}{\rho})\rVert^{2}_{F}$. Thus, the problem in (22) is transformed to
@@
Remark 4: After obtaining the learned consistent tensor $\mathcal{S}_{1}$ and view-specific tensor $\mathcal{S}_{2}$, we require to fuse them to achieve the final graph matrix $\mathbf{S}_{a}$. A natural fusion scheme is to average the accumulation of all the frontal slices of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, which is defined by
@@ IV Simulations
IV-A Experimental Settings
IV-A1 Datasets
We select six representative real-world multi-view datasets, a brief description on them is given below:
HW  [^1] https://cs.nyu.edu/$\sim$ roweis/data.html: It is composed of 2,000 data points for digits 0 and 9 from UCI machine learning repository and two public features are available.
Reuters  [^2] https://lig-membres.imag.fr/grimal/data.html: It contains 18,785 documents written in five different languages: French, Italian, English, German, and Spanish. We use 600 documents from 6 categories in our experiments.
BBCSport  [^3] http://mlg.ucd.ie/datasets/bbc.html: This dataset contains 544 documents sourced from the BBC Sport website, covering news from five different categories. In the experiments, two views with dimensions of 3,183 and 3,203 are selected.
3-sources  [^4] http://mlg.ucd.ie/datasets/3sources.html: This dataset includes 948 news articles covering 416 unique news stories. These articles are collected from three well-known online news sources. For our experiments, we use 169 articles that are classified into 6 categories across all three sources.
Yale  [^5] http://vision.ucsd.edu/content/yale-face-database: It consists of 165 gray-scale images of 15 individuals with different facial expressions and configurations. Followed by [39], 4096 dimensions intensity feature, 3,304 LBP feature and 6,750 dimensions Gabor feature are extracted as three multi-view features.
WebKB  [^6] http://www.webkb.org/webkb.html: This dataset consists of 203 web-pages of 4 classes. Each web-page is described by the content of the page, the anchor text of the hyper-link, and the text in its title.
@@ IV-A2 Compared Methods
MCGC [53]: This method proposes an effective disagreement cost function to learns the consensus graph.
@@ IV-B Experimental Results
Figure 2: Visualization of affinity matrix on BBCSport.
Figure 3: Parameter sensitivity results on four datasets.
Figure 4: Convergent behavior of CSTGL.
Figure 5: Comparison of running time of various multi-view methods.
IV-B1 Clustering Comparison
c) Generally, tensor-based multi-view methods outperform other non-tensor-based competing methods. For example, previous methods such as t-SVD-MSC and LTBPL demonstrate clustering performance that is only second to our proposed method on the HW dataset, achieving accuracies of $98.85\%$ and $99.9\%$, respectively. This superior performance can be attributed to the comprehensive consideration of high-order information by tensor decomposition. In contrast to these previous tensor-based methods, our proposed method, CSTGL, exhibits even better clustering performance. This can be largely attributed to its capability to capture the underlying structure of data on the Stiefel manifold, and the utilization of a newly designed tensor-based graph learning model that ensures both consistency and specificity, while effectively excluding the influence of noise.
TABLE IV: Ablation studies of CSTGL using different distance measures on various datasets, and the metric ACC is employed. (a) Euclidean distance, and (b) Stiefel manifold distance.
@@ IV-B1 Clustering Comparison
IV-B2 Ablation Studies
To validate the advantages of similarity measure on Stiefel manifold and tensor graph fusion, we conduct corresponding ablation studies. Specifically, when investigating the effects of using different distance measures, we alternate the similarity measure between Euclidean distance and Stiefel manifold distance while maintaining all other implementation procedures unchanged. When exploring the benefits of tensor graph fusion, we examine the effects of discarding the consistent tensor $\mathcal{S}_{1}$ and specific tensor $\mathcal{S}_{2}$, individually.
Table IV records the experimental results of CSTGL using different distance measures on various datasets. It is seen that for most tested datasets, the performance of CSTGL using Euclidean distance is inferior to that using Stiefel manifold distance. However, for the 3-sources dataset, CSTGL exhibits weaker performance when using the Stiefel manifold. This can be attributed to the nature of the 3-sources dataset, which is news-related, and the data may be sparse due to the limitation of news sources. In such a scenario, Euclidean distance, without considering the intrinsic structure of the data, could be more effective.
Table V shows the ablation studies of tensor graph fusion on the BBCSport, 3-sources, and WebKB datasets. From an overall perspective, both the consistent tensor graph and specific tensor graph contribute to the performance improvement of CSTGL. In particular, on the BBCSport and WebKB datasets, the consistent component contributes more significantly than the specific component, while on the 3-sources dataset, the specific component has a slightly greater contribution than the consistent component.
IV-B3 Visualization of Affinity Matrix
@@
IV-B4 Parameter Sensitivity Analysis
To measure the influence of parameter choices on the CSTGL’s performance, the parameter sensitivity analyses involving all the evaluation metrics on the BBCsport, HW, Yale and WebKB datasets are conducted, as shown in Fig. 3. By extensive experiments, we find that CSTGL is not sensitive to the selection of parameter $\alpha$. Consequently, we fix $\alpha=10$, and investigate the sensitivity induced by $\beta$ and $\gamma$. Overall, it is observed that the performance of CSTGL does not significantly vary with the choice of $\beta$. However, it does exhibit sensitivity to the choice of $\gamma$, particularly when $\beta$ is set to a small value. More specifically, when $\beta=10^{-3}$, a decrease in $\gamma$ leads to a deterioration in CSTGL’s performance. By referring to the objective function in (LABEL:eq014), we can see that the parameter $\gamma$ is used to regulate the view-specific tensor term. Therefore, we recommend using large $\gamma$ to emphasize the view’s specificity for desirable clustering performance.
IV-B5 Empirical Convergence Analysis
@@ V Conclusion
We propose a novel tensor-based graph learning framework, namely CSTGL. Our work makes two significant contributions in 1) the learning of neighbor graph, and 2) the model of graph fusion. Firstly, we learn the neighbor graph of each view based on using the Stiefel manifold distance, which is empirically verified to more effectively recover the intrinsic structure than using Euclidean distance. Secondly, after obtaining the learned neighbor graphs for all views, we formulate a tensor-based graph fusion model by assuming that the neighbor graph of each view is composed of a consistent graph, a view-specific graph, and a noise term. This model considers both consistency and specificity, while avoiding the issue of potential noise contamination in the learned specific information. Additionally, thanks to the t-SVD based tensor nuclear norm, CSTGL is capable of capturing high-order correlations among different views. Extensive experiments have demonstrated that CSTGL outperforms some SOTA multi-view methods.

### v3

@@ III-C Optimization
Let $e^{(v)}_{ij}=d_{v}-\mathrm{tr}(\mathbf{x}^{(v)}_{i}(\mathbf{x}^{(v)}_{j}))^{T}$, we can further reformulate the problem in (LABEL:eq017) separately for each $i$
@@
Given that the $l_{2,1}$-norm of a tensor is defined as the sum of the $l_{2}$-norms for each fiber along the third mode, thus we have $\lVert\mathbf{W}_{(3)}\rVert_{2,1}=\lVert\mathcal{W}\rVert_{2,1}$. Similarly, according to the definition of the $F$-norm of a tensor, it is straightforward to see $\lVert\mathbf{W}_{(3)}-(\mathbf{E}_{(3)}+\frac{\mathbf{Q}_{3_{(3)}}}{\rho})\rVert^{2}_{F}=\lVert\mathcal{W}-(\mathcal{E}+\frac{\mathcal{Q}_{3}}{\rho})\rVert^{2}_{F}$. Thus, the problem in (20) is transformed to
@@
After obtaining the learned consistent tensor $\mathcal{S}_{1}$ and view-specific tensor $\mathcal{S}_{2}$, we require to fuse them to achieve the final graph matrix $\mathbf{S}_{a}$. A natural fusion scheme is to average the accumulation of all the frontal slices of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, which is defined by
@@
Remark 4: Since our model is tensor-based, we cannot strictly guarantee that the final learned $\mathcal{S}_{1}$ satisfies the constraint $\mathcal{S}_{1}=\Phi(\mathbf{S},\cdots,\mathbf{S})$. This is an inherent limitation of the tensor learning process. However, we argue that this constraint does not conflict with the final learned result, as it represents an idealized prior assumption, whereas the learned result reflects a practical representation. Minor deviations from the ideal constraint are acceptable within the learning process. To support this, we have conducted heatmap experiments in the experimental section, which demonstrate that each frontal slice of the learned $\mathcal{S}_{1}$ exhibits only minor deviations, implying that $\mathcal{S}_{1}=\Phi(\mathbf{S},\cdots,\mathbf{S})$ approximately holds.
@@ IV Experiments
Figure 2: The Frobenius norm heatmaps between $\mathcal{S}_{1}^{(i)}$ and $\mathcal{S}_{1}^{(j)}$ on different datasets.
@@
IV-A Experimental Settings
IV-A1 Datasets
We select six representative real-world multi-view datasets, and a brief description on them is given below.
HW  [^1] https://cs.nyu.edu/$\sim$ roweis/data.html: It is composed of 2,000 data points for digits 0 and 9 from UCI machine learning repository. The digit images are obtained from two sources: MNIST and USPS Handwritten Digits. Accordingly, two views are constructed, with feature dimensions of 784 and 254, respectively.
Reuters  [^2] https://lig-membres.imag.fr/grimal/data.html: It contains 18,785 documents written in five different languages: French, Italian, English, German, and Spanish. Each view corresponds to a document representation in one of these languages. Following the widely employed selection strategy [28], we use 600 documents from 6 categories in our experiments.
BBCSport  [^3] http://mlg.ucd.ie/datasets/bbc.html: This dataset contains 544 documents sourced from the BBC Sport website, covering news from five different categories. Each view represents a different feature representation of the documents. In the experiments, two views with dimensions of 3,183 and 3,203 are selected.
3-sources  [^4] http://mlg.ucd.ie/datasets/3sources.html: This dataset includes 948 news articles covering 416 unique news stories, collected from three well-known online news sources. Each view corresponds to articles from one of these sources. Following the existing literature [28, 38], we use 169 articles that are classified into 6 categories across all three sources.
Yale  [^5] http://vision.ucsd.edu/content/yale-face-database: It consists of 165 gray-scale images of 15 individuals with different facial expressions and configurations. Followed by [37], three views are constructed based on different feature representations: a 4,096-dimensional intensity feature, a 3,304-dimensional LBP feature, and a 6,750-dimensional Gabor feature.
WebKB  [^6] http://www.webkb.org/webkb.html: This dataset consists of 203 web-pages of 4 classes. Each view represents a different textual aspect of a web page, including its main content, the anchor text of hyperlinks, and the text in its title.
@@ IV-A2 Compared Methods
MCGC [65]: This method proposes an effective disagreement cost function to learn the consensus graph.
@@
Figure 3: Visualization of affinity matrix on BBCSport.
Figure 4: Sensitivity analysis of $\alpha$ on different datasets, and $\beta=100$, $\gamma=10$.
@@ IV-B Experimental Results
IV-B1 Verification of Consistent Tensor Constraint
To demonstrate that each frontal slice of the learned $\mathcal{S}_{1}$ exhibits only minor deviations, we compare the Frobenius norm heatmaps of any two learned slices, $\mathbf{S}_{1}^{(i)}$ and $\mathbf{S}_{1}^{(j)}$, as shown in Fig. 2. It is clearly seen that on the tested datasets, every non-diagonal block displays very small values, indicating that the differences between any two learned slices are negligible. This empirical analysis confirms that the frontal matrices of the learned consistent tensor graph $\mathcal{S}_{1}$ are nearly identical, i.e., $\mathcal{S}_{1}=\Phi(\mathbf{S},\cdots,\mathbf{S})$ approximately holds.
TABLE IV: Ablation studies of CSTGL using different distance measures on various datasets, and the metric ACC is employed. (a) Euclidean distance, and (b) pseudo-Stiefel manifold distance.
@@ IV-B1 Verification of Consistent Tensor Constraint
TABLE VI: Ablation studies of tensor graph fusion on the Yale, Reuters and HW datasets. (a) retaining consistent tensor graph, (b) retaining view-specific tensor graph, and (c) retaining both consistent and view-specific tensor graphs.
Case Component Yale Reuters HW
Con. Spe. ACC NMI ARI Fscore ACC NMI ARI Fscore ACC NMI ARI Fscore
(a) 0.6545 0.7392 0.5530 0.5819 0.9950 0.9844 0.9880 0.9900 0.9995 0.9986 0.9989 0.9990
(b) 0.3273 0.3766 0.0990 0.1567 0.2133 0.0135 0.0014 0.1668 0.2110 0.0742 0.0353 0.1322
(c) 0.7697 0.7956 0.6381 0.6608 0.9967 0.9896 0.9920 0.9933 1.0000 1.0000 1.0000 1.0000
IV-B2 Clustering Comparison
c) Generally, tensor-based multi-view methods outperform other non-tensor-based competing methods. For example, previous methods such as t-SVD-MSC and LTBPL demonstrate clustering performance that is only second to our proposed method on the HW dataset, achieving accuracies of $98.85\%$ and $99.9\%$, respectively. This superior performance can be attributed to the comprehensive consideration of high-order information by tensor decomposition. In contrast to these previous tensor-based methods, our proposed method, CSTGL, exhibits even better clustering performance. This can be largely attributed to its capability to capture the underlying structure of high-dimensional data, and the utilization of a newly designed tensor-based graph learning model that ensures both consistency and specificity, while effectively excluding the influence of noise.
In the supplementary material, we present additional clustering comparison experiments with synthetic noise added to multi-view data. The results further demonstrate our model’s robustness against noise.
Figure 5: Parameter sensitivity results on six datasets, and $\alpha=10$.
Figure 6: Convergent behavior of CSTGL.
Figure 7: Comparison of running time of various multi-view methods.
IV-B3 Ablation Studies
To validate the advantages of similarity measure using pseudo-Stiefel manifold distance and tensor graph fusion, we conduct corresponding ablation studies. Specifically, when investigating the effects of using different distance measures, we alternate the similarity measure between Euclidean distance and pseudo-Stiefel manifold distance while maintaining all other implementation procedures unchanged. When exploring the benefits of tensor graph fusion, we examine the effects of discarding the consistent tensor $\mathcal{S}_{1}$ and specific tensor $\mathcal{S}_{2}$, individually.
Table IV records the experimental results of CSTGL using different distance measures on various datasets. It is seen that for most tested datasets, the performance of CSTGL using Euclidean distance is inferior to that using pseudo-Stiefel manifold distance. However, for the 3-sources dataset, CSTGL exhibits weaker performance when using the pseudo-Stiefel manifold distance. This can be attributed to the nature of the 3-sources dataset, which is news-related, and the data may be sparse due to the limitation of news sources. In such a scenario, Euclidean distance, without considering the intrinsic structure of the data, could be more effective.
Tables V and VI shows the ablation studies of tensor graph fusion on six datasets. From an overall perspective, both the consistent tensor graph and specific tensor graph contribute to the performance improvement of CSTGL. In particular, on the BBCSport, WebKB, Yale, Reuters, and HW datasets, the consistent component contributes more significantly than the specific component, while on the 3-sources dataset, the specific component has a slightly greater contribution than the consistent component.
IV-B4 Visualization of Affinity Matrix
@@
IV-B5 Parameter Sensitivity Analysis
To measure the influence of parameter choices on the CSTGL’s performance, the parameter sensitivity analyses involving all the evaluation metrics on six datasets are conducted. We first study the influence of $\alpha$ by fixing $\beta$ and $\gamma$. As shown in Fig. 4, the performance of CSTGL remains insensitive to $\alpha$ on the WebKB, BBCSport, 3-sources, and Yale datasets. However, a noticeable decline in performance is observed with large $\alpha$ on the Reuters and HW datasets. These observations indicate flexibility in choosing $\alpha$, with a preference of selecting small $\alpha$ for some real-world datasets.
We then investigate the sensitivity induced by $\beta$ and $\gamma$, and the experimental results are shown in Fig. 5. Overall, it is observed that the performance of CSTGL does not significantly vary with the choice of $\beta$. However, it does exhibit sensitivity to the choice of $\gamma$, particularly when $\beta$ is set to a small value. More specifically, when $\beta=10^{-3}$, a decrease in $\gamma$ leads to a deterioration in CSTGL’s performance. By referring to the objective function in (LABEL:eq014), we can see that the parameter $\gamma$ is used to regulate the view-specific tensor term. Therefore, we recommend using large $\gamma$ to emphasize the view’s specificity for desirable clustering performance.
IV-B6 Empirical Convergence Analysis
@@ V Conclusion
We propose a novel tensor-based graph learning framework, namely CSTGL. Our work makes two significant contributions in 1) the learning of neighbor graph, and 2) the model of graph fusion. Firstly, we learn the neighbor graph of each view based on using the pseudo-Stiefel manifold distance, which is empirically verified to more effectively recover the intrinsic structure than using Euclidean distance. Secondly, after obtaining the learned neighbor graphs for all views, we formulate a tensor-based graph fusion model by assuming that the neighbor graph of each view is composed of a consistent graph, a view-specific graph, and a noise term. This model considers both consistency and specificity, while avoiding the issue of potential noise contamination in the learned useful information. Additionally, thanks to the t-SVD based tensor nuclear norm, CSTGL is capable of capturing high-order correlations among different views. Extensive experiments have demonstrated that CSTGL outperforms some SOTA multi-view methods.